OpenStudy (anonymous):

The angle of elevation of a jet fighter from a point A on the ground is 60°. After a flight of 10 seconds, the angle of elevation changes to 30°. If the jet is flying at a speed of 432 km/hour, find the constant height at which the jet is flying.

OpenStudy (anonymous):

use \[\sqrt{3}=1.732\]

OpenStudy (anonymous):

my naswer which i know is wrong is coming 249.08 km

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@uri

OpenStudy (anonymous):

@Ashleyisakitty

OpenStudy (campbell_st):

ok... so the plane is flying 432 m/h so in 10 seconds it flies 432/(60*6) = 1.2 miles now draw a picture...  triangle BAJ is isosceles so JA = AB so you now have a right triangle  so you can use the sin ratio to find the height son(60) = h/1.2 hope it helps...